5. Fill in this table using these three ECG recordings. +Note the Q-wave is not always visible on this or all ECG recordings.ECG 1 entre 0 2:42 2:42.5 2:43 2:43.5 2:44 2:44.5 2:45 2:45.5 2:46 1.5 ECG 2 han har han mo 0.0 1.5 8 8.5 9 43 20 10.5 11 11.
Answer #1

Ecg 1 - RR interval it is the time between two qrs complex.heart rate can be calculated from this. In this ecg RR interval indicate the function of sinus node and it's intrinsic property

PR interval denotes the time from beginning of atrial depolarization to onset of ventricular depolarization

ST interval represents interval between normal depolarisation and repolarisation.

Heart rate is 300 divided by RR interval

300/7=42

Normal heart rate is 60-100 beats per min

In this it is Brady cardia

Ecg 2 - RR interval prolonged

PR interval normal

ST segment elevation

Heart rate - 300/12 =25/min

Indicates severe bradycardia

Ecg 3 RR interval appears normal

PR interval normal

ST segment elevation

Heart rate 300/5 =60

Normal heart rate